OpenStudy (anonymous):

Evaluate: [(40 + 5) − 32] ÷ 9 • 2

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

you can enter the same thing you wrote here into wolfram alpha, that link is an example. I just copied what you wrote and pasted it in. Check it out.

OpenStudy (anonymous):

that doesn't show her the steps though, how will she learn

OpenStudy (anonymous):

She'll learn about another resource that might be a great asset.

OpenStudy (anonymous):

[(40 + 5) − 32] ÷ 9 • 2 45-32÷ 9 • 2 13/9*2= 26/9

OpenStudy (anonymous):

giving her answers isn't gonna help her either. Have you heard of Pemda's @candeebabee ?

OpenStudy (anonymous):

pemdas

OpenStudy (anonymous):

lol @JenniferSmart1 people like that make me have a short giggle

OpenStudy (anonymous):

Okay, favorite mnemonic for Pemdas...

OpenStudy (anonymous):

Please excuse my dear aunt sallie

OpenStudy (anonymous):

@candeebabee do you have any questions?

OpenStudy (anonymous):

Please Eat My Dead Ant Salad

OpenStudy (anonymous):

@JenniferSmart1 I think she's been gone a while. Favorite mnemonic?

OpenStudy (anonymous):

sure LOL

OpenStudy (anonymous):

[(40 + 5) − 32] ÷ 9 • 2 = ? what is • at the?

OpenStudy (anonymous):

[(40 + 5) − 32] ÷ 9 • 2 [(45)-32] ÷ 9 • 2 [13]÷ 9 • 2 13/9 * 2 26/9 Which is your final answer